# Towards Intelligent Battery Management via A Five-Tier Digital Twin Framework

### v1

@@ Abstract
Battery management systems (BMSs) are critical to ensuring safety, efficiency, and longevity across electronics, transportation, and energy storage. However, with the rapid growth of lithium-ion batteries, conventional reactive BMS approaches face limitations in health prediction and advanced maintenance management, resulting in increased safety risks and economic costs. To address these challenges, we propose a five-tier digital twin framework for intelligent battery management. The framework spans geometric visualization, predictive modeling, prescriptive optimization, and autonomous operation, enabling full lifecycle optimization. In validation, an electrochemical model calibrated via Bayesian optimization achieved strong alignment with measured voltage and temperature, with Mean Absolute Percentage Errors (MAPE) below 1.57% and 0.39%. A Physics-Informed Neural Network (PINN) then combined data and simulations to predict State of Health (SOH), attaining MAPE under 3% with quantified uncertainty. This framework elevates BMSs into intelligent systems capable of proactive management and autonomous optimization, advancing safety and reliability in critical applications.
Index Terms:
Digital twin, Battery Management System (BMS), Bayesian optimization, Physics-informed Neural Network (PINN), State of Health (SOH) prediction
I Introduction
@@
Battery management systems (BMSs) are critical for ensuring the safety, efficiency, and longevity of batteries in Electric Vehicles (EVs) and grid energy storage. The indispensable role of BMSs is driven by the rapid expansion of energy storage application, expected to reach around 442 GWh globally by 2030 [1]. This accelerating deployment is not only increasing capacity demands but also introducing growing system complexity. For instance, EV battery packs contain thousands of cells operating under fluctuating loads, while grid-scale systems integrate heterogeneous batteries with diverse degradation patterns. These complexities demand BMSs capable of coordinating cell behavior and maintaining resilience under varying demand and supply conditions.
Conventional BMSs face significant challenges in managing dynamic operational conditions, with limited predictive capabilities and inadequate sensing that restricts the modeling of critical internal states. They typically rely on easily measurable signals such as current, terminal voltages, and surface temperatures, which provide limited visibility into latent states like lithium concentration gradients, internal temperature distributions, State of Charge (SOC), or State of Health (SOH) [2]. Without such detailed information, it becomes difficult to link sensor data to underlying degradation mechanisms or to optimize charge–discharge strategies.
Most existing BMS implementations remain rule-based designs, in which threshold-triggered protective actions provide essential short-term safety. This design arises because general BMS architectures rely on embedded Programmable Logic Controllers (PLCs) and a central Electronic Control Unit (ECU), which process only a limited set of measurable signals [3]. These controllers can activate local interlocks, shutdowns, or alarms to ensure immediate responses against incidents like short circuits or thermal runaway. While effective for mitigating immediate operational risks, this inherently reactive approach offers minimal predictive insights into long-term battery health and degradation mechanisms. To enable a more active, adaptive, and intelligent BMS, recent research has explored the integration of battery digital twins. A digital twin is a synchronized digital replica of the physical system that combines the asset itself, a high-fidelity virtual model, and bi-directional data flows to maintain alignment [4]. In batteries, digital twins often leverage electrochemical and thermal modeling, supported by remote monitoring and cloud-based computation, to deliver improved insights into internal states and predictive control [5].
Despite existing advancements in both traditional BMS and current digital twin solutions, there are several technical gaps mentioned above are still remain and need to be overcomed. First, existing research lacks a digital twin framework that tightly integrates high-fidelity multi-physics battery models with advanced AI techniques. Current battery digital twin approaches either rely on simplified physics or purely data-driven surrogates, which compromises both accuracy and generalization, making the virtual replica less interpretable and trustworthy. Second, the absence of a unified system-level framework means that the perception–prediction–control loop is handled in isolation, leading to fragmented insights and delayed responses. Consequently, current digital twin systems struggle to keep operations and virtual models aligned with the accuracy and speed needed for autonomous decision-making.
@@
To bridge these gaps, we propose a five-tier digital twin framework systematically evolving battery management from geometric modeling to autonomous control. The framework integrates real-time data assimilation at the descriptive tier, multi-physics simulation for accurate forecasting at the predictive tier, optimization algorithms for prescriptive control, and autonomous closed-loop operation at the highest tier, offering unprecedented predictive accuracy and operational intelligence. To realize these functions, the framework is validated through a co-simulation methodology. Electrochemical and thermal dynamics are modeled in physics-based simulators and synchronized within NVIDIA Omniverse, which provides a unified 3D virtual environment. Continuous-time prediction is supported by NVIDIA PhysicsNeMo to train physics-informed models that rapidly simulate real-world battery behaviors across diverse operating conditions. Reinforcement learning–based adaptation supports optimization under dynamic conditions, and closed-loop feasibility is tested through Hardware-in-the-Loop (HIL) validation. Different from conventional designs that treat monitoring, forecasting, and actuation as separate modules, this architecture establishes a unified digital intelligence pipeline evolving across the entire battery lifecycle, enabling interpretable modeling, reliable prediction, and intelligent control.
@@
  • To the best of our knowledge, this is the first article that systematically proposes and introduces a unified five-tier digital twin framework for intelligent battery management. Each tier is clearly defined to address the fundamental gaps in existing BMS approaches.
@@
  • To the best of our knowledge, this is the first framework that employs calibrated multi-physics modeling with PIML to achieve interpretable tracking of internal battery states and latent variables, which directly enhances the physical interpretability of battery management and supports proactive decision-making.
  • We further validate the effectiveness of our framework through a comprehensive case study on multi-physics model calibration and SOH prediction. Our approach achieves remarkable accuracy, reducing voltage and temperature simulation Mean Absolute Percentage Error (MAPE) to less than 1% through Bayesian optimization. Moreover, our physics-informed neural networks combined with uncertainty quantification methods achieve robust SOH predictions with a MAPE consistently below 3%, demonstrating substantial improvements over conventional battery management methods.
II PRELIMINARIES
@@
In this section, we first review BMS architectures and the limitations of conventional rule-based designs. We then introduce multi-physics modeling approaches for electrochemical, thermal, and degradation processes. Finally, we summarize key simulation platforms and validation methods that bridge virtual models with real-world operation.
II-A Overview of BMS
@@
BMSs are designed to monitor, protect, and optimize batteries. In practice, a BMS integrates subsystems for signal acquisition, central control, circuit protection, and communication with upper-level energy management systems [6]. Despite these capabilities, conventional implementations remain limited in intelligence, as they rely heavily on rule-based logic where protective actions such as disconnection, cooling, or cell balancing are triggered only after predefined safety thresholds are exceeded [7]. hile such designs mitigate immediate risks, they remain inherently reactive and inefficient, motivating the need for more advanced and predictive BMS solutions.
II-B Battery Multi-physics Modeling Techniques
@@
Battery multi-physics modeling sets up the foundation for estimating and predicting battery dynamic behavior based on electrochemical, chemical and mechanical principles. At the electrochemical level, broadly used models include the Single Particle Model (SPM) and the Doyle–Fuller–Newman (DFN) model are developed to model the temperature, cell voltage and other critical states [8]. The SPM simplifies battery dynamics by considering single representative particles, offering computational efficiency, while the DFN model provides detailed insights by simulating complex lithium-ion transport and electrochemical reactions across battery electrodes.
@@
Thermal modeling is a key component that characterizes heat generation, transfer, and dissipation during battery operation, typically formulated from energy conservation laws. Lumped thermal models [9] are widely adopted due to their simplicity and computational efficiency in estimating temperature dynamics, thereby supporting effective thermal management strategies essential for safety and reliability. Battery degradation modeling, on the other hand, focuses on long-term aging driven by mechanisms such as Solid Electrolyte Interphase (SEI) growth, lithium plating, and particle cracking. Advanced frameworks, such as those introduced by Wang et al. [10], explicitly couple these processes to capture the complex interactions that govern performance decay and material loss. These integrated multi-physics modeling plays an important role in accurately predicting battery lifetime and performance, emphasizing the importance and challenge of incorporating comprehensive multi-physics models within digital twin architectures. It is worth mentioning that both the thermal models and degradation models can be coupled within the multi-physics model, such as SPM and DFN, by introducing new source functions.
II-C Simulation Platforms and Validation Methods
@@
High-fidelity simulation platforms are a trustworthy way to validate battery digital twins, ensuring accurate representation of electrochemical, thermal, and mechanical behaviors under diverse conditions. Besides well-known tools like PyBaMM and COMSOL, other notable platforms include ANSYS Fluent for computational fluid dynamics and thermal analysis, and MATLAB/Simulink for dynamic system modeling and control strategy simulations. These platforms offer great support in analyzing and simulating battery operation, facilitating comprehensive case study evaluation and virtual testing without the necessity for costly physical prototype validations.
@@
HIL and Software-in-the-Loop (SIL) are two main complementary simulation paradigms that establish a bridge between virtual models and real-world systems [11]. In SIL, controller algorithms and software modules are embedded within a virtual simulation environment, allowing rapid prototyping, algorithm debugging, and iterative design without the need for physical hardware. Moreover, HIL extends this principle by incorporating real physical components into the simulation loop, thereby exposing algorithms to realistic operating conditions and hardware constraints. By combining these two approaches, digital twins benefit from a continuous validation pipeline that spans from early-stage software verification to hardware-level performance testing, laying the foundation for reliable predictive maintenance and operational optimization of batteries.
III Framework Design of five-tier digital twin for battery management
@@
In this section, we present the digital twin framework for battery management applications and BMS design. First, we introduce the overall framework design. Then, we articulate several potential applications of this digital twin in predictive and prescriptive battery analysis.
@@
Figure 1: Overview of the proposed five-tier digital twin intelligence framework, systematically integrating geometric modeling, descriptive analytics, predictive forecasting, prescriptive optimization, and autonomous control, enabling comprehensive lifecycle management and proactive decision-making for intelligent battery management.
III-A Architecture
Our five-tier digital twin framework is built upon three core technological modules that work synergistically and systematically: (i) a Virtual 3D Environment powered by NVIDIA Omniverse for geometric modeling and visualization, (ii) a multi-physics simulation engine for high-fidelity battery state estimation and prediction, and (iii) an AI engine leveraging NVIDIA PhysicsNeMo with PIML to tighten synchronization between the physical battery and its digital replica. Together, these modules enable progression across five distinct intelligence tiers, as shown in Figure 1. Each layer is constructed upon the foundation of the previous one, collectively enabling comprehensive battery lifecycle management.
@@ III-A Architecture
From bottom to top, Tier 1 (Geometric) builds high‑fidelity 3D representations of cells, modules, and packs, implemented with a Universal Scene Description (USD)‑based environment in NVIDIA Omniverse as the spatial backbone; Tier 2 (Descriptive) binds real‑time sensor streams to this geometry to create a live, data‑rich twin that visualizes temperature distributions, voltage/current profiles, and operating conditions; Tier 3 (Predictive) leverages PIML to forecast capacity‑degradation trajectories, estimate SOH and RUL, and quantify thermal‑runaway risk under diverse scenarios; Tier 4 (Prescriptive) translates these forecasts into optimal operating strategies, such as fast‑charging protocols, cell‑balancing schedules, and cooling‑system set‑points, while enforcing safety and operational constraints; and Tier 5 (Autonomous) enables the framework to achieve closed-loop, AI-driven control, where decisions are executed independently and models adapt dynamically through continual learning. This five-tier digital twin intelligence framework design is developed based on the following three key technologies:
III-A1 Virtual 3D Environment Module
@@
NVIDIA Omniverse serves as the foundational platform for constructing high-fidelity virtual representations of battery systems. USD files enable standardized geometric modeling spanning from individual cell components to complete battery system assemblies with many individual packs. Within this framework, this module is designed based on a USD-based 3D scene environment, which is completely programmable using a Python script to add or remove 3D objects and determine their spatial location. In our case, the scene graph captures the full battery hierarchy from cell to module to pack to represent a 20 kWh-level lithium-ion battery energy storage system. On top of this geometric structure of an energy storage system, the module adds a semantic schema that assigns meaning to each element. While the scene graph defines how cells, modules, and packs of the batteries are arranged in space, the semantic schema describes what each object represents and how it should be interpreted. For instance, a node can be marked as a temperature sensor, linked to the quantity it measures, and associated with a unit such as degrees Celsius. In this way, the schema ensures that the 3D scene is not just a geometric model but also a machine-readable map of physical roles and properties. To ensure that these semantic descriptions are usable across different scenarios in the digital twin, the module also fixes global coordinate frames, standardizes unit systems, records sensor locations, and assigns unique identifiers. These definitions make it possible to map live telemetry onto the 3D geometry without ambiguity and to exchange data consistently among different simulation and control tools.
@@
In addition, the 3D environment module defines live data mappings that connect telemetry streams and simulation outputs to the corresponding geometric entities in the scene. These mappings specify how data is transmitted and processed, including communication topics, update rates, and data types, while providing standardized input–output interfaces to connect with other tiers. The module also exposes read/write APIs for Tier 3 predictive services to present measured, estimated, and predicted results such as SOH and RUL in the 3D scene, while for Tier 4 controllers to visualize recommended setpoints. Figure 2 illustrates this integration within the Omniverse digital twin interface. This figure presents a 3D battery pack model augmented with live sensor dashboards and overlays of predicted SOH, RUL, and temperature variations results. Through this design, the shared 3D scene serves as the central hub where all information about the operational condition, real-time system information, and optimized control commands are integrated, ensuring consistency and coordination inside the battery digital twin.
@@
III-A2 Multi-physics Simulation Engine
Different from the preliminary modelling techniques that describe electrochemical, thermal, or degradation processes in isolation due to the mechanistic complexity, the simulation engine in the battery digital twin integrates these domains into a unified, co-evolving environment. It is designed to advance electrochemical, thermal, and mechanical solvers into a coupled and comprehensive model so that the digital twin reproduces battery dynamics in a physically consistent manner. For example, electrochemical reactions and ohmic resistance heat are identified as different heat generation sources that result in thermal dynamic variations. The rising temperature inside the cell then has an impact on the reaction kinetics and transport processes as described by the Arrhenius equation, reshaping voltage responses and accentuating spatial non-uniformities across cells. When mechanical effects are additionally considered, lithiation-induced swelling and assembly preload alter contact conditions and conductance pathways, thereby capturing the loss of interfacial integrity and localized resistance increases observed in real packs. By resolving these cause-and-effect loops, the multi-physics simulation engine can reproduce key phenomena such as hot-spot initiation during fast charging, thermal propagation across modules, and polarization shifts under transient loads.
This multi-physics simulation engine not only aligns with experimentally observed behaviours of actual battery systems but also produces high-quality synthetic datasets for the battery digital twin. It also conducts calibrated simulation campaigns in which operating profiles, ambient conditions, thermal management strategies, and aging states are systematically synchronised, generating multi-domain information in time series such as voltage, current, temperature fields, concentration distributions, and derived health indicators. Due to these outputs being computed in a coupled electrochemical–thermal–mechanical model, they are physically interpretable and can be used to deduce other critical latent states that are hard to observe using sensors. This enriches the battery health-related information, making it particularly valuable for supervised learning and benchmarking. Furthermore, the data reliability of this simulation engine is ensured through real-time calibration against experimental measurements to make the simulation responses match well with the observed responses. Additionally, physical consistency checks based on universal physical principles, such as energy conservation laws, are utilised to further reduce the mismatch between the digital twin battery and the physical battery.
@@ III-A2 Multi-physics Simulation Engine
III-A3 AI Engine Module
The AI engine provides powerful support for the intelligent control and accurate state monitoring of digital-twin batteries. Compared with conventional model-based approaches, AI offers more efficient and autonomous learning capabilities, enabling knowledge transfer and improved generalization. In our digital twin framework design, NVIDIA PhysicsNeMo serves as the foundation of the AI engine, enabling the battery digital twin to deliver real-time intelligent estimation, prediction and management while preserving physical consistency and interpretability. PhysicsNeMo follows a PIML paradigm that incorporates governing principles such as charge conservation, electrochemical kinetics, and thermodynamic constraints are embedded directly into neural architectures and training objectives, ensuring both estimations and predictions remain physically plausible and trustworthy rather than purely data-driven. PhysicsNeMo implements neural PDE solvers to accelerate electrochemical and thermal dynamics by several orders of magnitude compared to classical finite-element methods. Meanwhile, transformer-based or other sequence-learning models capture long-term dependencies that characterize degradation trajectories. Together, these methods provide robust forecasts of voltage response, temperature evolution, SOH, RUL, and associated uncertainty at battery cell, module, and pack levels.

### v2

# Towards Intelligent Systems for Battery Management: A Five-Tier Digital Twin Architecture

@@ Abstract
As digital twin technologies are increasingly incorporated into battery management systems to meet the growing need for transparent and lifecycle-aware operation, existing battery digital twins still suffer from fragmented operational processes and lack an architectural perspective to coordinate modeling, inference, and decision-making throughout the battery lifecycle. To this end, we develop a unified five-tier battery digital twin framework that integrates key functionalities into a coherent pipeline and facilitates a clearer architectural understanding of digital twins. The five-tier comprises geometric modeling, descriptive analytics, physics-informed prediction, prescriptive optimization, and autonomous control. In quantitative evaluation, the resulting architecture achieves high-fidelity multi-physics calibration with 0.92% voltage and 0.18% temperature prediction error, and provides state-of-health estimation with 1.09% MAPE and calibrated uncertainty. As the first battery digital twin system empowered by the NVIDIA ecosystem with physics-AI technologies, our proposed five-tier framework shifts battery management from reactive protection to an interpretable, predictive, and autonomous paradigm, paving the path to develop next-generation battery management and energy management systems.
[…]
@@
Battery Management Systems (BMSs) serve a critical function in ensuring the safety and performance of electric vehicles and energy storage systems. With the increasing scale of battery pack integration and the growing complexity of operating conditions, BMSs should not only ensure basic safety protection but also progressively expand their capabilities to meet the pressing demand for higher-level intelligent management. The indispensable role of BMSs is driven by the rapid expansion of energy storage applications, expected to reach around 442 GWh globally by 2030 [BNEF2023storage]. This accelerating deployment is increasing not only capacity demands but also introducing system complexity. For instance, EV battery packs contain thousands of cells operating under fluctuating loads, while grid-scale systems integrate heterogeneous batteries with diverse degradation patterns [cheng2016consumer]. These complexities demand BMSs capable of coordinating cell behavior and maintaining resilience under varying demand and supply conditions.
In practice, conventional BMSs remain limited in their sensing and prediction capabilities, primarily due to their reliance on a narrow range of sensor-accessible parameters. First, perception is limited because typical BMS architectures based on embedded PLCs and ECUs can process only a narrow set of measurable signals [muneer2025way], providing limited visibility into latent states like State of Charge (SOC), or State of Health (SOH) [cui2025towards]. This hardware limitation leads most existing BMSs to adopt rule-based designs with threshold-triggered protective actions for short-term safety. Second, predictive capability is limited because present BMSs mainly rely on simplified physics models that capture only short-term electrical responses while neglecting electrochemical degradation mechanisms. Consequently, they cannot simulate state trajectories under varying load and thermal conditions, making long-term health evolution prediction unattainable.
To overcome these limitations, researchers have introduced the concept of digital twins into BMSs, integrating multi-source data, multi-physics modeling, and advanced AI methods to establish a dynamic mapping between the physical and virtual domains. A digital twin is a synchronized digital replica of the physical system that combines the asset itself, a high-fidelity virtual model, and bi-directional data flows to maintain alignment [ali2021cognitive]. In the battery domain, digital twins typically leverage physics-based electrochemical models and multi-node thermal models to capture spatiotemporal distributions of temperature and lithium concentration. Because such high-fidelity simulations are computationally intensive, they are often executed on cloud or high-performance computing platforms, with the results synchronized back to the edge BMS for real-time decision support [xie2024dual]. As a result, current solutions tend to cover one or a few functional layers in isolation separately.
Despite recent progress, the fundamental gap in current battery digital twin frameworks is the absence of a unified, multi-tier architecture that organizes fragmented capabilities and workflows into a coherent whole. Building on this overarching architectural void, three specific gaps emerge. First, lacking synergy between data-driven models and physics-based models makes it difficult to simultaneously achieve high predictive accuracy and physical interpretability. In practice, many battery digital twin implementations rely on black-box data-driven approaches that capture correlations but ignore physical laws, while physics-based electrochemical or thermal models offer mechanistic insight but remain computationally intensive and struggle to represent degradation behaviors under practical cycling conditions. Second, most existing battery digital twins lack self-evolving capability. They still depend heavily on offline analysis, which prevents timely updates of the virtual model. As a result, the digital twin cannot continuously adapt to real-time operating data, leading to gradual divergence from the physical battery under dynamic conditions. Third, current systems lack decision validation capabilities. Although many digital twins can generate accurate predictions of battery behavior, they seldom verify how these predictions translate into effective operational decisions. Thus, the digital twin cannot ensure that its recommendations, such as charging or scheduling strategies, consistently improve real-world performance.
@@
To bridge these gaps, we propose a highly intelligent five-tier battery digital twin architecture empowered by the NVIDIA ecosystem, systematically organizing the full spectrum of digital twin capabilities. This battery digital twin integrates real-time data assimilation at the descriptive tier, multi-physics simulation for accurate forecasting at the predictive tier, optimization algorithms for prescriptive control, and autonomous closed-loop operation at the highest tier, forming a continuous pipeline from monitoring to intelligent actuation. Electrochemical and thermal dynamics are captured in physics-based simulators and synchronized within NVIDIA Omniverse, with NVIDIA SimReady assets streamlining standardized model construction, while NVIDIA PhysicsNeMo provides the physics-informed learning engine that accelerates simulation and forecasting. Unlike conventional designs that separate monitoring, forecasting, and control into isolated modules, our digital twin establishes an evolving digital intelligence across the entire battery lifecycle, enabling continuous synchronization and self-evolution between the physical and virtual systems, thereby delivering enhanced safety and reliability for next-generation EVs and energy storage applications.
@@
  • To the best of our knowledge, this is the first article that systematically proposes and introduces a unified five-tier digital twin system for intelligent battery management. Each tier is clearly defined to address the gaps in existing BMS approaches.
@@
  • The system uniquely employs calibrated multi-physics modeling with PIML to achieve interpretable tracking of internal battery states and latent variables, which directly enhances the physical interpretability of battery management and supports proactive decision-making.
  • We present the first battery digital twin system implemented within the NVIDIA ecosystem, leveraging Omniverse, SimReady assets, and PhysicsNeMo for end-to-end geometric modeling, multi-physics simulation, and physics-informed learning. Our implementation delivers high simulation fidelity, achieving low voltage and temperature prediction error, and produces robust SOH estimates with quantified uncertainty.
Preliminaries
@@
In this section, we first review the common BMS architectures. We then introduce multi-physics modeling approaches for electrochemical, thermal, and degradation processes. Finally, we summarize key simulation platforms and validation methods that bridge virtual models with real-world operation.
Overview of BMS
@@
BMSs are designed to monitor and protect batteries. In practice, a BMS integrates subsystems for signal acquisition, central control, circuit protection, and communication with upper-level energy management systems [gabbar2021review]. Despite these capabilities, conventional implementations remain limited in intelligence, as they rely heavily on rule-based logic where protective actions such as disconnection, cooling, or cell balancing are triggered only after predefined safety thresholds are exceeded [fioravanti2020predictive]. While such designs mitigate immediate risks, they remain inherently reactive and inefficient, motivating the need for more advanced and predictive BMS solutions.
Battery Multi-physics Modeling Techniques
@@
Battery physics modeling sets up the foundation for estimating and predicting battery dynamic behavior based on electrochemical, chemical and mechanical principles. At the electrochemical level, broadly used models include the Single Particle Model (SPM) and the Doyle–Fuller–Newman (DFN) model [doyle1993modeling]. The SPM simplifies battery dynamics by considering single representative particles, offering computational efficiency, while the DFN model provides detailed insights by simulating complex lithium-ion transport and electrochemical reactions across battery electrodes.
@@
Thermal modeling is a key component that characterizes heat generation, transfer, and dissipation during battery operation, typically formulated from energy conservation laws. Lumped thermal models [alkhedher2024electrochemical] are widely adopted due to their simplicity and computational efficiency in estimating temperature dynamics, thereby supporting effective thermal management strategies essential for safety and reliability. Battery degradation modeling, on the other hand, focuses on long-term aging driven by mechanisms such as solid electrolyte interphase growth, lithium plating, and particle cracking. Advanced modeling methods, such as those introduced by Wang et al. [wang2018review], explicitly couple these processes to capture the complex interactions that govern performance decay and material loss. These integrated multi-physics modeling plays an important role in accurately predicting battery lifetime and performance, emphasizing the importance and challenge of incorporating comprehensive multi-physics models within digital twin architectures. It is worth mentioning that both the thermal models and degradation models can be coupled within the multi-physics model by introducing new source functions.
Simulation Platforms and Validation Methods
@@
High-fidelity simulation platforms are a trustworthy way to validate battery digital twins, ensuring accurate representation of electrochemical, thermal, and mechanical behaviors under diverse conditions. Besides well-known tools like PyBaMM and COMSOL, other notable platforms include ANSYS Fluent for computational fluid dynamics and thermal analysis, and MATLAB/Simulink for dynamic system modeling and control strategy simulations. These platforms support analyzing and simulating battery operation, facilitating comprehensive virtual testing without the necessity for costly physical prototype validations.
@@
Hardware-in-the-Loop (HIL) and Software-in-the-Loop (SIL) are two main complementary simulation paradigms that establish a bridge between virtual models and real-world systems [bui2019advanced]. In SIL, controller algorithms and software modules are embedded within a virtual simulation environment, allowing rapid prototyping, algorithm debugging, and iterative design without the need for physical hardware. Moreover, HIL extends this principle by incorporating real physical components into the simulation loop, thereby exposing algorithms to realistic operating conditions and hardware constraints. By combining these two approaches, digital twins benefit from a continuous validation pipeline that spans from early-stage software verification to hardware-level performance testing, laying the foundation for reliable predictive maintenance and operational optimization of batteries.
System Design of the Digital Twin
@@
In this section, we present the intelligent five-tier digital twin system for battery management. First, we introduce the overall system architecture. Then, we articulate several potential applications of this digital twin in predictive and prescriptive battery analysis.
Architecture Overview
@@
Figure 1: Overview of the proposed five-tier digital twin intelligence system, integrating geometric modeling, descriptive analytics, predictive forecasting, prescriptive optimization, and autonomous control for intelligent battery management.
@@
The architecture of our intelligent five-tier digital twin system is shown in Figure 1. Each layer is constructed upon the foundation of the previous one within a unified intelligent system. From bottom to top, Tier 1 (Geometric) builds 3D representations of battery cells, modules, and packs, implemented in NVIDIA Omniverse as the spatial backbone, supported by NVIDIA SimReady assets that provide standardized, simulation-ready battery components; Tier 2 (Descriptive) binds real‑time sensor streams to this geometry to create a live, data‑rich twin that visualizes temperature distributions, voltage/current profiles, and operating conditions; Tier 3 (Predictive) leverages PIML to forecast capacity‑degradation trajectories, estimate SOH and RUL, and quantify thermal‑runaway risk under diverse scenarios; Tier 4 (Prescriptive) translates these forecasts into optimal operating strategies, such as fast‑charging protocols, cell‑balancing schedules, and cooling‑system set‑points, while enforcing safety and operational constraints; and Tier 5 (Autonomous) enables the system to achieve closed-loop, AI-driven control, where decisions are executed independently and models adapt through continual learning. NVIDIA Omniverse underpins Tier 1 and Tier 2 as the shared spatial and data plane, augmented by NVIDIA SimReady standardized battery assets, while PhysicsNeMo powers Tier 3 to Tier 5 for learning, prescription, and autonomy.
This intelligent system is enabled by three core modules spanning across tiers, with the overall organization illustrated in Figure 3: (i) a virtual 3D environment powered by NVIDIA Omniverse and NVIDIA SimReady for geometric modeling and visualization, (ii) a multi-physics simulation engine for high-fidelity battery state estimation and prediction, and (iii) an AI engine leveraging NVIDIA PhysicsNeMo with PIML to tighten synchronization between the physical battery and its digital replica.
Virtual 3D Environment Module
The virtual 3D environment underpins Tier 1 (Geometric) and Tier 2 (Descriptive) with NVIDIA Omniverse serving as the foundational platform for constructing high-fidelity virtual replicas of battery systems. As illustrated in Figure 3, this module forms the core of the Visualization Layer, where USD files enable standardized geometric modeling spanning from cell components to complete battery system assemblies with many packs. Within this system, this module is designed based on a USD-based 3D scene environment, which is completely programmable using a Python script to add or remove 3D objects and determine their spatial location. In our settings, the scene graph captures the full battery hierarchy from cell to module to pack to represent a 20 kWh-level lithium-ion battery energy storage system. To accelerate and standardize the construction of such virtual battery assemblies, we leverage NVIDIA SimReady assets, which provide preconfigured, simulation-ready models of cells, modules, and packs that integrate natively with the USD-based Omniverse environment.
@@
On top of this geometric structure of an energy storage system, the module adds a semantic schema that assigns meaning to each element. While the scene graph defines how cells, modules, and packs of the batteries are arranged in space, the semantic schema describes what each object represents and how it should be interpreted. For instance, a node can be marked as a temperature sensor, linked to the quantity it measures, and associated with a unit such as degrees Celsius. In this way, the schema ensures that the 3D scene is not just a geometric model but also a machine-readable map of physical roles and properties. To ensure that these semantic descriptions are usable across different scenarios in the digital twin, the module also fixes global coordinate frames, standardizes unit systems, records sensor locations, and assigns unique identifiers. These definitions make it possible to map live telemetry onto the 3D geometry without ambiguity and to exchange data consistently among different simulation and control tools.
@@
In addition, the 3D environment module defines live data mappings that connect telemetry streams and simulation outputs to the corresponding geometric entities. These mappings specify how data is transmitted and processed, including communication topics, update rates, and data types, while providing standardized input–output interfaces to connect with other tiers. The module also exposes read/write APIs for Tier 3 predictive services to present measured, estimated, and predicted results such as SOH and RUL in the 3D scene, while for Tier 4 controllers to visualize recommended setpoints. Figure 2 illustrates this integration within the Omniverse digital twin interface. This figure presents a 3D battery pack model augmented with live sensor dashboards and overlays of predicted SOH, RUL, and temperature variations results. Through this design, the shared 3D scene serves as the central hub where all information about the operational condition, real-time system information, and optimized control commands are integrated, ensuring consistency and coordination inside the battery digital twin.
@@
Multi-physics Simulation Engine
The multi-physics simulation engine underpins Tier 3 (Predictive) by enabling high-fidelity forecasting of battery dynamics through coupled electrochemical, thermal, and mechanical modeling. As shown in Figure 3, this module constitutes the core of the Model Layer and Simulation Engine, which serve as the analytical backbone of the overall system. Simulation outputs are streamed to Omniverse for live visualization and are used as supervised and physics constrained signals for model training and inference in PhysicsNeMo.
The simulation engine integrates these physics domains into a unified, co-evolving environment, ensuring that the digital twin reproduces battery behavior in a physically consistent manner. For example, electrochemical reactions and ohmic resistance heat are identified as different heat generation sources that result in thermal dynamic variations. The rising temperature inside the cell then impacts the reaction kinetics and transport processes as described by the Arrhenius equation [laidler1984development], reshaping voltage responses and accentuating spatial non-uniformities across cells. By resolving these cause-and-effect loops, the multi-physics simulation engine can capture critical battery behaviors such as hot-spot formation, heat spread, and performance shifts under varying loads.
This multi-physics simulation engine not only aligns with experimentally observed behaviours of actual battery systems but also produces high-quality synthetic datasets for the battery digital twin. It also conducts calibrated simulation campaigns in which operating profiles, ambient conditions, thermal management strategies, and aging states are systematically synchronised, generating multi-domain information in time series such as voltage, current, temperature fields, concentration distributions, and derived health indicators. Due to these outputs being computed in a coupled physics model, they are physically interpretable and can be used to deduce other critical latent states that are hard to observe using sensors. This enriches the battery health-related information, making it particularly valuable for supervised learning and benchmarking. Furthermore, the data reliability of this simulation engine is ensured through real-time calibration against experimental measurements to make the simulation responses match well with the observed responses. Additionally, physical consistency checks based on universal physical principles, such as energy conservation laws, are utilised to further reduce the mismatch between the digital twin battery and the physical battery.
@@
AI Engine Module
The AI engine supports Tier 4 (Prescriptive) and Tier 5 (Autonomous) by providing intelligent optimization and adaptive control capabilities for the battery digital twin. As illustrated in Figure 3, it represents the intelligence core of the overall system within the Machine Learning Platform, where advanced AI methods complement physics-based modeling. In our digital twin system design, NVIDIA PhysicsNeMo serves as the foundation of the AI engine, enabling the battery digital twin to deliver real-time intelligent estimation, prediction and management while preserving physical consistency and interpretability. PhysicsNeMo follows a PIML paradigm that incorporates governing principles such as charge conservation, electrochemical kinetics, and thermodynamic constraints are embedded directly into neural architectures and training objectives, ensuring both estimations and predictions remain physically plausible and trustworthy rather than purely data-driven. PhysicsNeMo implements neural PDE solvers to accelerate electrochemical and thermal dynamics by several orders of magnitude compared to classical finite-element methods. Meanwhile, transformer-based or other sequence-learning models capture long-term dependencies that characterize degradation trajectories. Together, these methods provide robust forecasts of voltage response, temperature evolution, SOH, RUL, and associated uncertainty at battery cell, module, and pack levels.